OpenStudy (anonymous):

Urgent!!!! medal Cody and Earl started week 1 of their garden with 20 tomatoes. Each week they eat 10 tomatoes and sell all that remain, but they grow 3 times as many as they sell. How many tomatoes will they have at the start of week 5? a. 420 c. 150 b. 1570 d. 1620

OpenStudy (anonymous):

@phi help again

OpenStudy (anonymous):

It's A,420

OpenStudy (anonymous):

If they start out with 20 tomatoes and eat ten per week they are giving away the other 10, yes?

OpenStudy (here_to_help15):

20-10 = 10 Week 1 10 * 3 = 30 30 - 10 = 20 Week 2 20 * 3 = 60 60 - 10 = 50 Week 3 50 * 3 = 150 150 -10 = 140 Week 4 140 * 3 = 420 So at the start of week 5 they have 420

OpenStudy (anonymous):

@Here_to_Help15 knows whats up!

OpenStudy (here_to_help15):

yup : )

OpenStudy (anonymous):

thanks

OpenStudy (here_to_help15):

No prob guess this i answered instead of asking you a question : P

OpenStudy (anonymous):

:D yeah and thank u for that